OpenStudy (anonymous):

the x-intercept of a line is (-2 1/2,0) and the y-intercept is (0,5) what is the equation of the line in slope-intercept form?

OpenStudy (campbell_st):

well find the slope of the line  the general form is y = mx + b m = slope of the line b = y-intercept (which you are given, y = 5) just substitute the values

OpenStudy (anonymous):

i dont get where 5 goes ? into y=mx+b?

OpenStudy (campbell_st):

that is b

OpenStudy (campbell_st):

so the equation is y = mx + 5 that's given... you just need to slope

OpenStudy (anonymous):

y=mx+5?

OpenStudy (campbell_st):

thats the basic equation go to the diagram and calculate the slope...

OpenStudy (anonymous):

i just dont get it

OpenStudy (anonymous):

y=2x+5?

OpenStudy (campbell_st):

yep... that's it the rise = 5, run = 2.5 slope m = 5/2.5 or m = 2 so the equation is y = 2x + 5 the value of b, the y-intercept, is where the line cuts the vertical axis